# A Conflict-Aware Resource Management Framework for the Computing Continuum

### v1

@@ I Introduction
The growing complexity and heterogeneity of modern distributed systems have accelerated the shift toward the computing continuum, a paradigm that seamlessly integrates cloud, fog, and edge resources to provide low-latency, context-aware computing capabilities. Driven by the rapid proliferation of connected devices and data-intensive applications (e.g., smart cities to real-time industrial automation), this integrated computing model enables optimizing computation across various tiers and ensuring adaptability and efficiency under varying workloads and resource constraints [pujol2023distributed, pujol2023edge].
Orchestration frameworks such as Kubernetes have become essential for deploying and scaling services across computing continuum infrastructures. Mechanisms for managing dynamic resource conflicts in multi-agent, multi-objective optimization scenarios are still not well researched [truyen2019comprehensive]. As deployments expand into increasingly heterogeneous and latency-sensitive environments, traditional resource management approaches, which are often static or purely reactive, fall short of meeting modern demands[awad2022slo]. More specifically, computing infrastructures typically run a wide range of services, each with distinct performance requirements and resource usage patterns. These environments are inherently multi-agent [soumplis2024performance, zhang2019heteroedge, chen2025agentflow], where numerous components act concurrently to manage resources for different services. Resource conflicts arise when multiple agents issue contradictory actions. In this context, a conflict refers to incompatible resource specifications issued by distinct agents that cannot be simultaneously satisfied without degrading service performance or violating Service Level Objectives (SLOs). For example, during peak load, we might need to increase CPU allocation for critical services, while at the same time, another policy might aim to reduce infrastructure or device energy usage due to the high operating costs at a specific time. These competing goals can lead to resource conflicts, where multiple agents issue contradictory actions. One agent may attempt to reduce CPU usage to reduce costs, while another simultaneously requests more resources to maintain performance. These conflicting actions can create instability and cause repeated reconfigurations, degrade both system efficiency and service availability [tuli2023neurosurgeon, naderializadeh2021resource].
@@ II Related Work
The computing continuum aims to create a dynamic and adaptive computing environment that spans between cloud and edge computing [murturi2023learning, murturi2021decentralized]. A critical point emphasized in this research is the inadequacy of the reactive management system for computing continuum systems due to the unique demands.
Li et al.[li2025adaptive] introduce a graph-augmented multi-agent reinforcement learning system where local agents optimize resources in their neighborhoods, while a global orchestrator handles system-wide coordination. This collaborative MARL architecture effectively resolves competing goals among agents across heterogeneous infrastructure layers. However, the proposed approach is given as a high-level architecture without evaluation. Several important contributions in this field utilize DL and RL techniques to address various challenges. Most of these efforts primarily focus on resource allocation and management; however, they often neglect the crucial aspect of reactive interventions when conflicts arise between agents’ actions in the system. A notable example is the DRL model proposed by Zeng et al. [zeng2019resource]. This model offers a model-free solution designed to automatically predict user resource consumption patterns by analyzing historical interaction data. While this model presents a significant advantage through its model-free nature, it is constrained by its dependency on the specific environment in which it is deployed. Naderializadeh et al. [naderializadeh2021resource] suggest a DRL distributed model, which runs across multiple agents. This paper describes a more complex deep learning model that depends on the communication between the distributed agents. An important aspect of this paper is that the agents are independent and may make decisions simultaneously while being unaware of the other agent’s decisions. An interesting DRL implementation is presented by Liu et al.[liu2017hierarchical], which uses Q-Learning to derive the optimal choices when facing specific situations. In contrast to existing research, which primarily focuses on resource allocation optimization or workload prediction in isolated domains of the computing continuum, this work introduces a conflict-aware orchestration mechanism that addresses coordination challenges arising from user-defined resource management specifications. The proposed approach integrates a feedback-driven, self-adaptive mechanism that continuously learns from system performance and previous optimization outcomes.
@@ III-D Deep Reinforcement Learning (DRL)
The DRL system is invoked when runtime performance deteriorates following an applied specification or when logical conflicts remain unresolved by the watcher. In such cases, it adaptively selects corrective actions such as scaling, migration, or resource reallocation. The DRL system operates as a service hosting a neural network model trained to make optimization decisions for targeted deployments. Furthermore, the DRL system relies on real-time metrics such as CPU utilization, memory usage, and latency which are retrieved through the Metrics API. Two complementary mechanisms form the basis of the system. The general-purpose Meta DRL Model captures common optimization patterns and supports system-wide decision-making. When this meta model yields suboptimal outcomes for specific services, an Adaptive DRL mechanism via Instance-Based Learning retrains specialized models using feedback data. These service-specific models evolve over time, improving the precision of optimization actions. The following subsections describe the meta-level model for general optimization and the adaptive mechanism that enables reinforcement learning at the instance level [guo2020learning].

### v2

@@ I Introduction
The growing complexity and heterogeneity of modern distributed systems have accelerated the shift toward the computing continuum, a paradigm that seamlessly integrates cloud, fog, and edge resources to provide low-latency, context-aware computing capabilities. Driven by the rapid proliferation of connected devices and data-intensive applications (e.g., smart cities to real-time industrial automation), this integrated computing model enables optimizing computation across various tiers and ensuring adaptability and efficiency under varying workloads and resource constraints [9, 10].
Orchestration frameworks such as Kubernetes have become essential for deploying and scaling services across computing continuum infrastructures. Mechanisms for managing dynamic resource conflicts in multi-agent, multi-objective optimization scenarios are still not well researched [12]. As deployments expand into increasingly heterogeneous and latency-sensitive environments, traditional resource management approaches, which are often static or purely reactive, fall short of meeting modern demands[1]. More specifically, computing infrastructures typically run a wide range of services, each with distinct performance requirements and resource usage patterns. These environments are inherently multi-agent [11, 14, 2], where numerous components act concurrently to manage resources for different services. Resource conflicts arise when multiple agents issue contradictory actions. In this context, a conflict refers to incompatible resource specifications issued by distinct agents that cannot be simultaneously satisfied without degrading service performance or violating Service Level Objectives (SLOs). For example, during peak load, we might need to increase CPU allocation for critical services, while at the same time, another policy might aim to reduce infrastructure or device energy usage due to the high operating costs at a specific time. These competing goals can lead to resource conflicts, where multiple agents issue contradictory actions. One agent may attempt to reduce CPU usage to reduce costs, while another simultaneously requests more resources to maintain performance. These conflicting actions can create instability and cause repeated reconfigurations, degrade both system efficiency and service availability [8].
@@ II Related Work
The computing continuum aims to create a dynamic and adaptive computing environment that spans between cloud and edge computing [6, 7]. A critical point emphasized in this research is the inadequacy of the reactive management system for computing continuum systems due to the unique demands.
Li et al.[4] introduce a graph-augmented multi-agent reinforcement learning system where local agents optimize resources in their neighborhoods, while a global orchestrator handles system-wide coordination. This collaborative MARL architecture effectively resolves competing goals among agents across heterogeneous infrastructure layers. However, the proposed approach is given as a high-level architecture without evaluation. Several important contributions in this field utilize DL and RL techniques to address various challenges. Most of these efforts primarily focus on resource allocation and management; however, they often neglect the crucial aspect of reactive interventions when conflicts arise between agents’ actions in the system. A notable example is the DRL model proposed by Zeng et al. [13]. This model offers a model-free solution designed to automatically predict user resource consumption patterns by analyzing historical interaction data. While this model presents a significant advantage through its model-free nature, it is constrained by its dependency on the specific environment in which it is deployed. Naderializadeh et al. [8] suggest a DRL distributed model, which runs across multiple agents. This paper describes a more complex deep learning model that depends on the communication between the distributed agents. An important aspect of this paper is that the agents are independent and may make decisions simultaneously while being unaware of the other agent’s decisions. An interesting DRL implementation is presented by Liu et al.[5], which uses Q-Learning to derive the optimal choices when facing specific situations. In contrast to existing research, which primarily focuses on resource allocation optimization or workload prediction in isolated domains of the computing continuum, this work introduces a conflict-aware orchestration mechanism that addresses coordination challenges arising from user-defined resource management specifications. The proposed approach integrates a feedback-driven, self-adaptive mechanism that continuously learns from system performance and previous optimization outcomes.
@@ III-D Deep Reinforcement Learning (DRL)
The DRL system is invoked when runtime performance deteriorates following an applied specification or when logical conflicts remain unresolved by the watcher. In such cases, it adaptively selects corrective actions such as scaling, migration, or resource reallocation. The DRL system operates as a service hosting a neural network model trained to make optimization decisions for targeted deployments. Furthermore, the DRL system relies on real-time metrics such as CPU utilization, memory usage, and latency which are retrieved through the Metrics API. Two complementary mechanisms form the basis of the system. The general-purpose Meta DRL Model captures common optimization patterns and supports system-wide decision-making. When this meta model yields suboptimal outcomes for specific services, an Adaptive DRL mechanism via Instance-Based Learning retrains specialized models using feedback data. These service-specific models evolve over time, improving the precision of optimization actions. The following subsections describe the meta-level model for general optimization and the adaptive mechanism that enables reinforcement learning at the instance level [3].